# A Survey of Text-to-SQL in the Era of LLMs: Where are we, and where are we going?

### v5

@@ II-C Text-to-SQL Task Challenges
  • Lexical Ambiguity: This occurs when a single word has multiple meanings. For example, the word “bat” can refer to an animal or a baseball bat or the action of swinging.
@@ II-D3 PLM-based Stage
The introduction of PLMs like BERT [41] and T5 [42] in 2018 led to significant advancements in Text-to-SQL methods based on PLMs [43, 7, 44], achieving competitive performance on various benchmarks (see Figure 2(b)-③). At this stage, PLMs-based Text-to-SQL models trained on large corpora have greatly enhanced natural language understanding, resolving approximately 80% of cases in the Spider dataset [45]. However, accuracy drops to about 50% on the extra hard cases of Spider [46]. In addition, these models still face challenges in handling complex schemas.
@@ II-D4 LLM-based Stage
where k denotes additional information or domain-specific knowledge related to nl or db. $P$ is a prompt function that transforms the input (nl, db, k) into a suitable textual prompt for the LLMs. An well-designed $P$ can effectively guide the LLMs to perform the Text-to-SQL task more accurately.
@@ III Language Model-powered Text-to-SQL Overview
  • Schema Linking: This module identifies the most relevant tables and columns from Text-to-SQL (Section IV-A).
@@ IV Pre-Processing Strategies for Text-to-SQL
The pre-processing step is crucial in the Text-to-SQL translation process, as it identifies relevant tables and columns (i.e., Schema Linking) and retrieves necessary database contents or cell values (i.e., DB Content Retrieval) to support sql query generation What’s more, it enriches context by incorporating domain-specific knowledge (i.e., Additional Information Acquisition), which can improve the understanding of the query context and correct errors to prevent their propagation.
@@ IV-A Schema Linking
We categorize existing schema linking strategies into three groups based on their characteristics: 1) string matching-based schema linking, 2) neural network-based schema linking, and 3) in-context learning-based for schema linking.
@@ IV-A2 Neural Network-based Schema Linking
DAE [88] frames schema linking as a sequential tagging problem, using a two-stage anonymization model to capture semantic relationships between schema and nl. SLSQL [53] annotates schema linking information in Spider dataset [45], enabling a systematic, data-driven study. RESDSQL [7] introduces a ranking-enhanced encoding framework for schema linking, using a cross-encoder to prioritize tables and columns based on classification probabilities. FinSQL [50] uses a parallel cross-encoder to retrieve relevant schema elements, significantly reducing linking time.
@@ IV-B1 String Matching-based Methods
IRNet [84] uses n-grams, treating text between quotes as cell values. BRIDGE [83] advances this with an anchor text matching technique that automatically extracts cell values from the nl. Using heuristics, it calculates the maximum sequence match to define matching boundaries, excluding irrelevant substrings and adjusts thresholds for accuracy.
@@ IV-B2 Neural Network-based Methods
TABERT [91] uses a method called database content snapshots to encode relevant database content for snl query, employing attention mechanisms to manage information across cell value representations in different rows. Another approach leverages graph relationships to represent database content. IRNet [84] uses the knowledge graph ConceptNet [92] to find and link relevant cell values, assigning types based on exact or partial matches. RAT-SQL [80] further enhances structural reasoning by modeling the relationship between cell values and the nl query, identifying column-value relationships where the query value is part of the column’s candidate cell value.
@@ IV-B3 Index Strategy for Database Content Retrieval
CHESS [59] uses a Locality-sensitive Hashing [93] for approximate nearest neighbor searches, indexing unique cell values to quickly find the top matches related to the nl query. This approach speeds up the process of comparing edit distances and semantic embeddings. CodeS [49] employs a coarse-to-fine matching strategy. It uses BM25 [94] to build an index for coarse-grained searches, identifying candidate values, which are then refined by applying the Longest Common Substring algorithm [95] to assess similarity with the nl, thereby pinpointing the most relevant cell values.
@@ V-B Decoding Strategy
Decoding is a crucial step in Text-to-SQL translation, transforming encoder-generated representations into sql. An effective decoding strategy ensures that the generated sql are not only syntactically correct but also semantically aligned with the nl queries, while optimizing sql execution efficiency. Figure 7 introduces several key decoding strategies.
@@ VI-B Output Consistency
DAIL-SQL [6] integrates self-consistency, achieving a $0.4\%$ performance improvement over configurations without it. To reduce LLM randomness, Fin-SQL [50] generates $n$ candidate sql queries in parallel, clusters them based on keyword consistency, and selects a query from the largest cluster. The self-consistency strategy enhances LLM output diversity by increasing the temperature and selecting the final result through majority voting. However, recent studies [118] indicate that relying on a single model may still yield limited diversity. To overcome this limitation, PET-SQL [62] introduces a cross-consistency strategy, in which multiple LLMs generate sql at lower temperatures and vote based on execution results.
@@ VI-C Execution-Guided Strategies
To reflect human behavior in writing complex sql queries, CHESS [59] provides LLMs with the database schema, question, candidate sql queries, and their execution results. CHESS starts with a draft query and refines it based on execution feedback, adjusting for syntax errors as needed. CodeS [49] on the other hand, generates a complete sql statements through beam search, producing four sql candidates and selecting the first executable one as the final result.
@@ VII-A An Overview of Text-to-SQL Benchmarks
Cross-Domain Text-to-SQL Datasets. After the development of early single-domain datasets, the Text-to-SQL field shifted toward cross-domain datasets to test systems’ generalization across diverse sql queries and databases. WikiSQL [110] was the first cross-domain dataset, drawing tables from Wikipedia across various domains. Subsequently, Spider [45] was introduced, containing more complex relational databases with multiple tables. Recently, the BIRD [52] has further advanced complexity by including sql functions and operations absent in Spider, providing a greater challenge for Text-to-SQL.
Multi-Turn Text-to-SQL Datasets. With advancements in Text-to-SQL, multi-turn datasets have been developed to support interactive dialogues. SParC [128] is a cross-domain and multi-turn dataset with about $4.3K$ nl questions, totaling over $12K$ (nl, sql) pairs, each nl questions requiring contextual understanding across turns. CoSQL [129], collected using a Wizard-of-Oz setup, includes over $30K$ turns and introduces additional challenges like unanswerable questions, further testing context comprehension.
@@ VII-B In-depth Analysis of Existing Text-to-SQL Datasets
Regarding Query Complexity, datasets like FIBEN [133] and SEDE [142] feature sql queries with multiple tables and aggregate functions, mirroring complexities in real-world financial domains and Stack Exchange sites. Recent datasets also emphasize Scalar Functions and Mathematical Computations, adding structural challenges.
@@ VIII-A Evaluation Metrics
$EM=\frac{\sum_{i=1}^{N}\mathbbm{1}(\bigwedge_{C_{k}\in\mathbb{C}}Y_{i}^{C_{k}}=\hat{Y}_{i}^{C_{k}})}{N}.\vskip-5.0pt$
@@
$\small VES=\frac{\sum_{i=1}^{N}\mathbbm{1}(V_{i}=\hat{V}_{i})\cdot\mathbf{R}(Y_{i},\hat{Y_{i}})}{N},~\mathbf{R}(Y_{i},\hat{Y_{i}})=\sqrt{\frac{\mathbf{E}(Y_{i})}{\mathbf{E}(\hat{Y}_{i})}},\vskip-4.62497pt$
where $\mathbf{R}(\cdot)$ measures the relative execution efficiency of the predicted sql query compared to the ground-truth sql query, eliminating uncertainties due to machine status. $\mathbf{E}(\cdot)$ measures the efficiency of specific sql query, which can be refer to execution time, memory usage and more.
Query Variance Testing (QVT) [46]. This metric measures the robustness of an Text-to-SQL system in handling variations in nl queries. For a given sql query $Y_{i}$, there are often multiple corresponding nl queries, represented as pairs {($Q_{1}$, $Y_{i}$), ($Q_{2}$, $Y_{i}$), …, ($Q_{m}$, $Y_{i}$)}. The QVT metric is calculated as:
@@ VIII-C A Taxonomy for Text-to-SQL Errors Analysis
Error analysis involves examining model errors to identify limitations and guide corrective actions for improved performance. In this section, we first review the existing Text-to-SQL error taxonomy. We then propose design principles and introduce a two-level Text-to-SQL errors taxonomy.
Existing Taxonomies for Text-to-SQL Errors Analysis. Recent Text-to-SQL research [33, 151, 152, 59, 5, 153] has increasingly incorporated error analysis, proposing various error taxonomies. Ning et al. [152] introduced a detailed error taxonomy based on two dimensions: (1) Syntactic dimension identifies specific sql parts where errors occur, organized by keywords such as WHERE and JOIN. (2) Semantic dimension indicates misinterpretations of the natural language description, such as errors in understanding table names. SQL-PaLM [153] categorizes errors into five types: (1) Schema Linking, irrelevant or missing table/column selection; (2) Database Content, misinterpreting data values; (3) Knowledge Evidence, failing to utilize external hints; (4) Reasoning, lacking intermediate logical steps; and (5) Syntax, invalid SQL format. NL2SQL-BUGs [33] focuses on the analysis of semantic errors, organizing them into 9 main categories and 31 subcategories. It further proposes a new benchmark for evaluating models’ error detection capabilities, advancing automated error analysis in the Text-to-SQL.
@@
Taxonomy Principles for Text-to-SQL Errors Analysis. Current error taxonomies in Text-to-SQL are often specific to particular datasets, limiting their general applicability. To address these issues, a standardized and effective taxonomy is essential. We propose the following principles [154] to guide the development of an Text-to-SQL error taxonomy:
@@ X Limitations and Open Problems
Interpreting Text-to-SQL Solutions. Understanding the reasoning behind a Text-to-SQL model’s performance enhances confidence in its reliability. Explainable AI techniques [156, 157], such as surrogate models [158] and saliency maps [159], aim to reveal model decisions. However, their effectiveness in Text-to-SQL contexts, especially with combined LLM and PLM, remains an open question. In addition, multi-agent LLM frameworks [160] improve reliability by splitting Text-to-SQL into specialized sub-tasks. Although this approach improves robustness, coordinating agents to ensure consistent and optimized performance remains a major challenge.

### v6

@@ II-C Text-to-SQL Task Challenges
  • Lexical Ambiguity: This occurs when a single word has multiple meanings. For example, the word “bat” can refer to an animal, or a baseball bat, or the action of swinging.
@@ II-D3 PLM-based Stage
The introduction of PLMs like BERT [41] and T5 [42] in 2018 led to significant advancements in Text-to-SQL methods based on PLMs [43, 7, 44], achieving competitive performance on various benchmarks (see Figure 2(b)-③). At this stage, PLM-based Text-to-SQL models trained on large corpora have greatly enhanced natural language understanding, resolving approximately 80% of cases in the Spider dataset [45]. However, accuracy drops to about 50% on the extra hard cases of Spider [46]. In addition, these models still face challenges in handling complex schemas.
@@ II-D4 LLM-based Stage
where k denotes additional information or domain-specific knowledge related to nl or db. $P$ is a prompt function that transforms the input (nl, db, k) into a suitable textual prompt for the LLMs. A well-designed $P$ can effectively guide the LLMs to perform the Text-to-SQL task more accurately.
@@ III Language Model-powered Text-to-SQL Overview
  • Schema Linking: This module identifies the most relevant tables and columns for Text-to-SQL (Section IV-A).
@@ IV Pre-Processing Strategies for Text-to-SQL
The pre-processing step is crucial in the Text-to-SQL translation process, as it identifies relevant tables and columns (i.e., Schema Linking) and retrieves necessary database contents or cell values (i.e., DB Content Retrieval) to support sql query generation. What’s more, it enriches context by incorporating domain-specific knowledge (i.e., Additional Information Acquisition), which can improve the understanding of the query context and correct errors to prevent their propagation.
@@ IV-A Schema Linking
We categorize existing schema linking strategies into three groups based on their characteristics: 1) string matching-based schema linking, 2) neural network-based schema linking, and 3) in-context learning for schema linking.
@@ IV-A2 Neural Network-based Schema Linking
DAE [88] frames schema linking as a sequential tagging problem, using a two-stage anonymization model to capture semantic relationships between schema and nl. SLSQL [53] annotates schema linking information in the Spider dataset [45], enabling a systematic, data-driven study. RESDSQL [7] introduces a ranking-enhanced encoding framework for schema linking, using a cross-encoder to prioritize tables and columns based on classification probabilities. FinSQL [50] uses a parallel cross-encoder to retrieve relevant schema elements, significantly reducing linking time.
@@ IV-B1 String Matching-based Methods
IRNet [84] uses n-grams, treating text between quotes as cell values. BRIDGE [83] advances this with an anchor text matching technique that automatically extracts cell values from the nl. Using heuristics, it calculates the maximum sequence match to define matching boundaries, excluding irrelevant substrings, and adjusts thresholds for accuracy.
@@ IV-B2 Neural Network-based Methods
TABERT [91] uses a method called database content snapshots to encode relevant database content for the nl query, employing attention mechanisms to manage information across cell value representations in different rows. Another approach leverages graph relationships to represent database content. IRNet [84] uses the knowledge graph ConceptNet [92] to find and link relevant cell values, assigning types based on exact or partial matches. RAT-SQL [80] further enhances structural reasoning by modeling the relationship between cell values and the nl query, identifying column-value relationships where the query value is part of the column’s candidate cell value.
@@ IV-B3 Index Strategy for Database Content Retrieval
CHESS [59] uses Locality-sensitive Hashing [93] for approximate nearest neighbor searches, indexing unique cell values to quickly find the top matches related to the nl query. This approach speeds up the process of comparing edit distances and semantic embeddings. CodeS [49] employs a coarse-to-fine matching strategy. It uses BM25 [94] to build an index for coarse-grained searches, identifying candidate values, which are then refined by applying the Longest Common Substring algorithm [95] to assess similarity with the nl, thereby pinpointing the most relevant cell values.
@@ V-B Decoding Strategy
Decoding is a crucial step in Text-to-SQL translation, transforming encoder-generated representations into sql. An effective decoding strategy ensures that the generated sql queries are not only syntactically correct but also semantically aligned with the nl queries, while optimizing sql execution efficiency. Figure 7 introduces several key decoding strategies.
@@ VI-B Output Consistency
DAIL-SQL [6] integrates self-consistency, achieving a $0.4\%$ performance improvement over configurations without it. To reduce LLM randomness, FinSQL [50] generates $n$ candidate sql queries in parallel, clusters them based on keyword consistency, and selects a query from the largest cluster. The self-consistency strategy enhances LLM output diversity by increasing the temperature and selecting the final result through majority voting. However, recent studies [118] indicate that relying on a single model may still yield limited diversity. To overcome this limitation, PET-SQL [62] introduces a cross-consistency strategy, in which multiple LLMs generate sql at lower temperatures and vote based on execution results.
@@ VI-C Execution-Guided Strategies
To reflect human behavior in writing complex sql queries, CHESS [59] provides LLMs with the database schema, question, candidate sql queries, and their execution results. CHESS starts with a draft query and refines it based on execution feedback, adjusting for syntax errors as needed. CodeS [49], on the other hand, generates a complete sql statements through beam search, producing four sql candidates and selecting the first executable one as the final result.
@@ VII-A An Overview of Text-to-SQL Benchmarks
Cross-Domain Text-to-SQL Datasets. After the development of early single-domain datasets, the Text-to-SQL field shifted toward cross-domain datasets to test systems’ generalization across diverse sql queries and databases. WikiSQL [110] was the first cross-domain dataset, drawing tables from Wikipedia across various domains. Subsequently, Spider [45] was introduced, containing more complex relational databases with multiple tables. Recently, BIRD [52] has further advanced complexity by including sql functions and operations absent in Spider, providing a greater challenge for Text-to-SQL.
Multi-Turn Text-to-SQL Datasets. With advancements in Text-to-SQL, multi-turn datasets have been developed to support interactive dialogues. SParC [128] is a cross-domain and multi-turn dataset with about $4.3K$ nl questions, totaling over $12K$ (nl, sql) pairs, each nl question requiring contextual understanding across turns. CoSQL [129], collected using a Wizard-of-Oz setup, includes over $30K$ turns and introduces additional challenges like unanswerable questions, further testing context comprehension.
@@ VII-B In-depth Analysis of Existing Text-to-SQL Datasets
Regarding Query Complexity, datasets like FIBEN [133] and SEDE [142] feature sql queries with multiple tables and aggregate functions, mirroring complexities in the real-world financial domains and Stack Exchange sites. Recent datasets also emphasize Scalar Functions and Mathematical Computations, adding structural challenges.
@@ VIII-A Evaluation Metrics
$EM=\frac{\sum_{i=1}^{N}\mathbbm{1}(\bigwedge_{C_{k}\in\mathbb{C}}Y_{i}^{C_{k}}=\hat{Y}_{i}^{C_{k}})}{N}.$
@@
$\small VES=\frac{\sum_{i=1}^{N}\mathbbm{1}(V_{i}=\hat{V}_{i})\cdot\mathbf{R}(Y_{i},\hat{Y_{i}})}{N},~\mathbf{R}(Y_{i},\hat{Y_{i}})=\sqrt{\frac{\mathbf{E}(Y_{i})}{\mathbf{E}(\hat{Y}_{i})}},$
where $\mathbf{R}(\cdot)$ measures the relative execution efficiency of the predicted sql query compared to the ground-truth sql query, eliminating uncertainties due to machine status. $\mathbf{E}(\cdot)$ measures the efficiency of specific sql query, which can refer to execution time, memory usage and more.
Query Variance Testing (QVT) [46]. This metric measures the robustness of a Text-to-SQL system in handling variations in nl queries. For a given sql query $Y_{i}$, there are often multiple corresponding nl queries, represented as pairs {($Q_{1}$, $Y_{i}$), ($Q_{2}$, $Y_{i}$), …, ($Q_{m}$, $Y_{i}$)}. The QVT metric is calculated as:
@@ VIII-C A Taxonomy for Text-to-SQL Errors Analysis
Error analysis involves examining model errors to identify limitations and guide corrective actions for improved performance. In this section, we first review the existing Text-to-SQL error taxonomies. We then propose design principles and introduce a two-level Text-to-SQL errors taxonomy.
Existing Taxonomies for Text-to-SQL Errors Analysis. Recent Text-to-SQL research [33, 151, 152, 59, 5, 153] has increasingly incorporated error analysis, proposing various error taxonomies. Ning et al. [152] introduced a detailed error taxonomy based on two dimensions: (1) Syntactic dimension identifies specific sql parts where errors occur, organized by keywords such as WHERE and JOIN. (2) Semantic dimension indicates misinterpretations of the natural language description, such as errors in understanding table names. SQL-PaLM [153] categorizes errors into five types: (1) Schema Linking, irrelevant or missing table/column selection; (2) Database Content, misinterpreting data values; (3) Knowledge Evidence, failing to utilize external hints; (4) Reasoning, lacking intermediate logical steps; and (5) Syntax, invalid SQL format. NL2SQL-BUGs [33] focuses on the analysis of semantic errors, organizing them into 9 main categories and 31 subcategories. It further proposes a new benchmark for evaluating models’ error detection capabilities, advancing automated error analysis in Text-to-SQL.
@@
Taxonomy Principles for Text-to-SQL Errors Analysis. Current error taxonomies in Text-to-SQL are often specific to particular datasets, limiting their general applicability. To address these issues, a standardized and effective taxonomy is essential. We propose the following principles [154] to guide the development of a Text-to-SQL error taxonomy:
@@ X Limitations and Open Problems
Interpreting Text-to-SQL Solutions. Understanding the reasoning behind a Text-to-SQL model’s performance enhances confidence in its reliability. Explainable AI techniques [156, 157], such as surrogate models [158] and saliency maps [159], aim to reveal model decisions. However, their effectiveness in Text-to-SQL contexts, especially with combined LLMs and PLMs, remains an open question. In addition, multi-agent LLM frameworks [160] improve reliability by splitting Text-to-SQL into specialized sub-tasks. Although this approach improves robustness, coordinating agents to ensure consistent and optimized performance remains a major challenge.
@@ XI Conclusion
Acknowledgments
This paper was supported by the National Key R&D Program of China (2023YFB4503600); the NSF of China (62402409, 62525202, 62232009, 62436010, and 62441230); the Guangdong Basic and Applied Basic Research Foundation (2023A1515110545); the Guangzhou Basic and Applied Basic Research Foundation (2025A04J3935); the Guangzhou-HKUST(GZ) Joint Funding Program (2025A03J3714); and the Guangdong provincial project (2023CX10X008).